A plate of area 10 cm^2 is to be electroplated with copper (density 9000 kg/m^3) to a thickness of 10 micrometres on both sides, using a cell of 12 V. Calculate the energy spent by the cell in the process of deposition. If this energy is used to heat 100 g of water, calculate the rise in the temperature of the water. ECE of copper = 3 × 10^-7 kg/C and specific heat capacity of water = 4200 J/kg-K.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. A= 10 cm2× 10^–4cm^2 t = 10m = 10 × 10^–6 Volume = A(2t) = 10 × 10^–4 × 2 × 10 × 10^–6 = 2 × 10^2 × 10^–10 = 2 × 10^–8 m^3 Mass = 2 × 10^–8 × 9000 = 18 × 10^–5 kg W = Z × C ⇒ 18 × 10^–5 = 3 × 10^–7 × C ⇒ q = 18 * 10^-5/3 * 10^-7 = 6 * 10^2 V = W/q = ⇒ W = Vq = 12 × 6 × 10^2 = 76 × 10^2 = 7.6 KJ